# Finite phase coherence time of a quantum field created by an ideal Bose gas

### v5

@@ Abstract
A quantitative quantum field approach for a very weakly interacting, dilute Bose gas is presented. Within the presented model, which assumes the constraint of particle number conservation at constant average energy in the canonical ensemble, both coherent oscillations, as well as decay times of quantum coherence for a quantum field created by the atomic cloud of a Bose-Einstein condensate, are modeled simultaneously by a unique complex time variable and two different characteristic frequencies for the oscillation and decoherence of the field. Within the present theory, it is illustrated that the occurrence of coherence and a macroscopic ground state population has its origin in finite coherence times of the ensemble of quantum particles in the Bose gas, which - in contrast to the incoherent interactions between the different particles - leads to the preparation of a thermodynamically stable many-body quantum state with coherent superpositions of discrete and quantized condensate and non-condensate atom number states at constant total atom number.
@@
As a fundamental property of quantum mechanics, coherence demonstrates the wave nature of quantum particles by the ability of the particles to interfere when their wavelength becomes of the order of their average distance below the critical temperature [2]. In classical physics, when a particle is moved from point A to a second point B, the particle has a well-defined position and velocity at each instant of time during the cycle from A to B, and indeed this time is well-defined by the path length between the two positions and the average velocity of the traveling particle [3]. Quantum mechanically, however, the dynamics of particles are more complicated, because of uncertainty and localization of position and velocity for particles in an external confinement, thus in the setup as described above, the quantum particle may only be located at two different positions and therefore exhibit no probability to be located in the space between the two points A and B. As a consequence of quantum effects from quantization of the corresponding single-particle wave functions, if the particle tunnels from point A to point B, there remains a time interval during which it is quantum mechanically not well-defined whether the particle is located at position A or B, so that the quantum state of the particle may thus be in a coherent superposition of a state with localization in point A and a state with localization in point B, with a well-defined phase relation between these two states which allows to quantify the oscillations of the transition probability for the particle to move from point A to B, and vice versa. The time scale during which these two states are quantum mechanically correlated and the particle’s wave function can interfere and coherently oscillate is called coherence time [4].
@@
Calculations and analyses of coherence times typically rely on the knowledge of the spectrum of the given quantum system. In a Bose-Einstein condensate below the critical temperature, most of the particles in the Bose gas search to populate the ground state of the system, so that a macroscopic occupation of the ground state mode with a typically centered spatial distribution occurs below the critical temperature for Bose-Einstein condensation with the characteristics of a coherent matter wave [5], since all the particles’ wave functions tend to interfere at the center of the external confinement. This happens because each of the particles can in principle tunnel from the ground state to a (multimode) excited state of the single-particle spectrum, because of the large energy uncertainty below the critical temperature. As a consequence of such single-particle coherence from tunneling as described above, each particle is thus in general in a superposition of a state with the localization characteristics of the single-particle ground state, and one or multiple of the excited single-particle states. The large coherence time achieved this way is a precursor for Bose-Einstein condensation, where the coherence time $t$ is required to be sufficiently large to allow the different single-particle wave functions to interfere constructively and coherently - i. e. in particular larger than the oscillation time of particles among the different single-particle quantum states a particle in the Bose gas may coherently populate.
Within the framework of quantum mechanics, quantum processes obey time-reversal symmetry due to well-defined phase relations, but only on the time scale of coherence, or more precisely, as long as the quantum system is not coupled to an external thermal environment. Thus, it is the coherence time that defines a natural unit for the forward propagation of time in a quantum system like a Bose-Einstein condensate which adjusts its direction of time due to decoherence by the maximization of entropy. However, in the singularity of a Bose-Einstein condensate, nothing ad hoc guarantees that the laws of standard thermodynamics, like the entropy maximization principle, still hold. This is not least, because the process of decoherence is not necessarily dominant in the parameter range of Bose-Einstein condensation, which typically leads to long-range order with only one characteristic many-body quantum state and large coherence times, in particular for ideally closed setups [6]. In such case, the single-particle entropy may remain constant, or even decrease in the evolution of the condensation process if standard thermodynamic properties break down in the singularity.
Figure 1: (color online) Figures show the spectrum of fugacity values for a Bose-Einstein condensate with $N=1000$ particles in a harmonic trap with trap frequencies $\omega_{x},\omega_{y}=2\pi\times 42.0$ Hz, and $\omega_{z}=2\pi\times 120.0$ Hz at different temperatures, $T=5.0{\rm~{}nK}$ (upper left), $T=10.0{\rm~{}nK}$ (upper right), $T=22.5{\rm~{}nK}$ (lower left) and $T=35.0{\rm~{}nK}$ (lower right). Below the critical temperature $T_{c}=26.9$ nK, there’s only one Boltzmann equilibrium state (Bose-Einstein condensation), which corresponds to ${\rm Re}(z)\rightarrow 1$ ($t\rightarrow 0$). In this limit, unconnected parts of the spectrum are strongly coupled. Whereas the fugacity ring spectrum is gapless below the critical temperature, it obeys a gap and tends to get continuously close above $T_{c}$ ($T=35.0{\rm~{}nK}$) with a large range of possible and coupled equilibrium states, obeying the same absolute value of the fugacity tends to one.
So far, no quantitative estimates for the coherence time of a quantum particle have been derived from first principles in physical units within standard field theories for Bose-Einstein condensation which make use of particle number conservation in the canonical ensemble. In the following analysis, the coherence time is quantified for a quantum particle in a nearly ideal Bose-Einstein condensate confined in a harmonic trap, starting from a number-conserving quantum field theory for Bose-Einstein condensation [8, 9], assuming sufficiently dilute atomic gases. In this context, it is shown that gauging the quantum field is a necessary and sufficient condition for spontaneous symmetry breaking of time reversal and phase gauge symmetry of the quantum field below the critical temperature, provided that the equilibrium state of the gas follows a Boltzmann equilibrium, which maximizes the entropy of the system in the present model. Time reversal symmetry and the corresponding phase gauge symmetry of a quantum particle in a Bose-Einstein condensate turn out to be spontaneously broken due to fluctuations of the corresponding entropy which does not rely on coherent particle interactions in this parameter regime, coupling symmetric and asymmetric parts of the underlying fugacity spectrum starting for temperatures at the critical temperature. The fact that the coherence time and the maximum entropy at the Boltzmann equilibrium are finite below the critical temperature can be considered as a signature for spontaneous symmetry breaking and therefore particularly indicates the quantum mechanical character of the Boltzmann equilibrium statistics in the present parameter range.
@@
In the present approach, complex time is formally modeled as a scaled two-dimensional complex variable - which arises intrinsically from the constraint of particle number conservation. Standard time reversal symmetry is recovered in the classical limit of large temperatures. To connect the mathematical framework of Bose-Einstein condensates at thermal equilibrium to the physics of coherence and oscillation times of a single particle in a Bose gas at finite temperature, the concept of imaginary time turns out to be a suitable mathematical tool to describe both the quantitative scaling of coherence times as well as their uncertainty and broken gauge symmetry aspects induced by many-particle fluctuations below the critical temperature [7, 10, 8]. Whereas the definition of (purely) imaginary time often lacks a clear physical interpretation in other frameworks, such as the calculation of the Hawking radiation in black holes [10], or the definition and calculation of a Eukledian time in special relativity [11], in the present framework, the formal splitting of the time variable into two variables derived from scaling the standard time parameter $t$ with the corresponding oscillation and decay frequency which arise from the underlying complex fugacity spectrum, imaginary time can be interpreted as the time scale during which a quantum particle in the gas relaxes towards an equilibrium quantum state imposing the assumption of a Boltzmann equilibrium (i. e. coupling elementary quantum states to the external environment). As shown in the sequel of this work, this is because the real part of complex time effectively converges to zero, whereas the imaginary part defines the relaxation time towards the Boltzmann equilibrium with maximal entropy at the given temperature. In contrast to (purely) imaginary time that defines the decay characteristics in the vertical direction of the complex plane, from the conservation of particle number and average energy in the framework of the canonical ensemble, the particles in the gas simultaneously exhibit both finite coherence times as well as oscillations between different regions A and B in space (e. g. spatial domains of different single-particle wave functions) below the critical temperature also at the Boltzmann equilibrium, an effect that can be energetically understood as the coupling of symmetric and asymmetric parts of the underlying complex fugacity spectrum. Thus, the foundation of coherent particle coupling among different condensate and non-condensate atom number states in a Bose-Einstein condensate as described in the present model is due to a process related to elementary time evolution which defines a finite coherence time that is larger or equal to the complex time variable which defines the (upper bound of a) time scale for a particle to interact with its (thermal) environment. Therefore, each of the particles is coherently distributed among different single-particle quantum states, which leads to a coherent field distribution with different possible condensate and non-condensate number states at a constant total particle number in the Bose gas, which finally defines a macroscopic ground state population from coherent wave interference at a stable Boltzmann equilibrium of maximum entropy below the critical temperature.
@@
where $\psi_{0}$ represents the spatially integrated condensate field, and $\psi_{\perp}$ the non-condensate field in the framework of solid numbers in the standard complex plane [8]. In the absence of external perturbations (below the critical temperature), and assuming isotropic relative spatial phase dependence, the total field $\psi=0$ vanishes so that the total symmetry of the system remains preserved. This is because the two condensate and non-condensate field components can be decomposed into uncorrelated single field components. We assume a spatially uniform distribution of the relative phase between the different single-particle wave functions. In contrast, as will be shown in the sequel of the present numerical analysis, below the critical temperature, where different atom number states get effectively coherently coupled because of finite coherence times, the resulting quantum states are of the form $|\mu\rangle=\sum_{\bf{k}}c_{\bf{k}}|\{N_{\bf{k}}\}\rangle$ which leads to the fact that $\psi_{0}=-\psi_{\perp}\neq 0$ called spontaneously broken gauge symmetry.
@@
hence defines a relation between the absolute temperature and the coherent dynamics of a quantum particle in the presence of the $N-1$ other particles in the Bose gas by its absolute value $|t|=\beta\hbar$. The time parameter $t$ can be interpreted as the absolute value of time in units of unit time of the equilibrated quantum gas (which is defined by the lowest single-particle energy of a quantum particle in the external confinement, since all values lower than the single-particle ground state energy are effectively zero, because of the finite energy uncertainty), or equivalently as the single-particle entropy in units of the Boltzmann constant $k_{B}$.
Figure 2: (color online) Figure shows $10^{5}$ realizations of the imaginary parts (upper panel - labeled $t_{1}$) and real parts (lower figure - labeled $t_{2}$) of complex time parameterized by $t=t_{1}+it_{2}$ as defined in Eq. (4) from random Markov sampling for a non-isotropic trap geometry $\omega_{x},\omega_{y}=2\pi\times 42.0$ Hz and $\omega_{z}=120.0$ Hz at temperature $T=5.0$ nK. The observation that $t_{2}$ is non-zero at thermal equilibrium indicates the coupling of the equilibrium state to quantum states with $t_{2}\neq 0$ (non-classical correlations).
Figure 2: (color online) Figure shows $10^{5}$ realizations of the imaginary parts (upper panel - labeled $t_{1}$) and real parts (lower figure - labeled $t_{2}$) of complex time parameterized by $t=t_{1}+it_{2}$ as defined in Eq. (4) from random Markov sampling for a non-isotropic trap geometry $\omega_{x},\omega_{y}=2\pi\times 42.0$ Hz and $\omega_{z}=120.0$ Hz at temperature $T=5.0$ nK. The observation that $t_{2}$ is non-zero at thermal equilibrium indicates the coupling of the equilibrium state to quantum states with $t_{2}\neq 0$ (non-classical correlations).
@@
As a formal ansatz, in Eq. (2), one may use that $\mu t$ is not a real, but a complex number, that is $\mu t\in\mathbb{C}$. Hence, in the framework of this quantum field theory as presented, the definition of scaled time consists of a real part $t_{1}$ which describes the coherent phase evolution of the quantum particle at equilibrium (inverse oscillation frequency in the stationary state), corresponding to the ring of the fugacity spectrum in Fig. 1, whereas $t_{2}$, the imaginary part (of $\mu t$) describes decay processes of the particle in the atomic cloud at or close to equilibrium (compare linear part of the fugacity spectrum in Fig. 1), that is
$\mu t=t_{1}+it_{2},$ (3)
which can be quantified by solving the equation for the constraint of particle number conservation numerically, see Eq. (5). Please note that the introduction of a two-dimensional (scaled) time is not entirely necessary for further calculations and results. In principle, the quantities $t_{1}$ and $t_{2}$ can also be represented by corresponding rates $\Gamma_{1}$ and $\Gamma_{2}$ weighted with a one-dimensional time variable $t$ [17]. However, the notation of two different time variables $t_{1}$ and $t_{2}$ (that is labeling the oscillation rate $\Gamma_{1}$ and the decay rate $\Gamma_{2}$, respectively) can bear some formal advantages, in particular in the framework of string theory [14]. Therefore, we will keep this notion for the analysis of this study. The absolute value of the parameter $t$ can also be interpreted as the total entropy of a quantum particle in units of the Boltzmann constant $k_{B}$, after a single particle excitation in the presence of the atomic background.
Above the critical temperature, the real part of the fugacity spectrum shows a gap between the ring of constant absolute time, indicating that the symmetry of the quantum field can in principle not be broken without externally induced quantum fluctuations of energy and the corresponding entropy. Decreasing temperature by external cooling to decrease the gap between the outer ring and the inner linear (symmetry breaking) part of the fugacity spectrum finally leads to a gapless fugacity spectrum. However, due to the uncertainty of complex time, there is also a finite coupling probability between symmetric and asymmetric parts of the fugacity spectrum, in the parameter range, where the fugacity spectrum obeys a gap between the two principally unconnected spectra close above the critical temperature, which leads to spontaneous symmetry breaking. Thus, a well-defined gauge of the quantum field (i. e. particle number conservation that leads to $\psi=0$) assuming Boltzmann equilibrium turns out to be a necessary and sufficient condition for spontaneous symmetry breaking of the quantum field [13], since the coherence time is finite starting from temperatures approaching the critical temperature for Bose-Einstein condensation (see Fig. 3).
@@
Figure 3: (color online) Figure shows the (scaled) coherence time $t_{2}$ as a function of temperature $T$ for a particle in the trap geometry $\omega_{x},\omega_{y}=2\pi\times 42.0$ Hz and $\omega_{z}=120.0$ Hz. For intermediate values of temperature, the coherence time is around $(300-450)~{}\mu$s. Above the critical temperature, it approaches a value of around $200~{}\mu$s.
@@
$\Delta N=\sum_{j\neq 0}z^{j}(\mu)\left[\prod_{k\neq x,y,z}\frac{1}{1-e^{-j\beta\hbar\omega_{k}}}-1\right]+\mathcal{O}(a).$ (5)
In Eq. (5), $\Delta N=(N-N_{0})$ is the difference between $N$ the total number of particles, and $N_{0}$ the number of condensate particles. Further, $\mu$ denotes the chemical potential of the Bose gas, and $\omega_{k}$ the trap frequency in mode direction $k=x,y,z$. For numerical calculations, the single-particle ground state energy (level) $0.5\hbar(\omega_{x}+\omega_{y}+\omega_{z})$ can be set to (effectively) zero.
Equation (4) does not necessarily describe a coherent forward or backward propagation in time, but only the probability for a transition (or new event) from a state with (the characteristics) of a quantum state with entropy $t_{A}$ to a state with entropy $t_{B}$, that is there is a non-vanishing probability that a transition with $t_{A}\geq t_{B}$ is accepted. Most likely, however, a transition is only accepted, if the absolute value $t_{B}\geq t_{A}$, which means that the particle tends to populate states with large entropy (and coherence time) in the Markov evolution process of the present quantum theory. In the present framework, the representation of a complex time arises naturally from Eq. (5), rather than from purely mathematical definition [14] and the system is allowed to exchange energy with the environment. In this setup, i. e. assuming the thermodynamical constraints of the canonical ensemble, coupling of unconnected parts of the fugacity spectrum isn’t possible without coherently coupled single-particle states as indicated by the finite (scaled) coherence time $t_{2}$. Since the variable $t$ can also be interpreted as the thermodynamic entropy in units of the Boltzmann constant, imposing Eq. (4) shows that the equilibrium state is a thermal equilibrium many-body quantum state of maximum entropy.
@@
For quantitative numerical modeling and analysis, the Bose gas is assumed to be dilute and obey a constant number of particles at finite temperature in a harmonic trapping potential to numerically calculate the coherence time (real part of complex time as defined in Eq. (2)) of a quantum particle in the Bose gas in SI units, which corresponds to the range of few hundred microseconds at the given parameter space. Interactions in the Bose gas are assumed to be negligible in the limit of a nearly ideal gas, where the s-wave scattering length $a$ of the particles in the gas effectively tends to zero [16]. From Eq. (2), we learn that the Bose gas then obeys a stable (Boltzmann) equilibrium for absolute values of complex time in the limit, where $t_{1}\rightarrow 0$ and $|t_{2}|\ll\hbar\beta$ (which corresponds to a small decoherence rate as compared to the upper bound $\hbar\beta$ for equilibration in the Bose gas). This indicates that the main decoherence mechanism in the Bose gas is thermalization.
@@
Applying only the constraint of particle number conservation as defined by the conservation equation in Eq. (5), numerical sampling of complex time leads to a distribution of typical time scales for $t_{1}$ (scaled oscillation time) and $t_{2}$ (scaled coherence time), shown in Fig. 2. From the sampling of complex time, it is observed that $t_{1}$, the real part of complex time, is distributed around zero (with a vanishing width of approximately $0.03~{}\mu s$) in the given parameter range (as defined in Fig. 2 - middle panel). The real part $t_{1}$ defines the average oscillation frequency in the fugacity spectrum of Fig. 2, which means that the quantum particle (with constant particle number at finite temperature) quickly couples to excited single particle quantum states, related to $t_{1}\sim 0\pm 0.03~{}\mu s$ in the form of a standing wave. The distribution of the time scale $t_{2}$, the directional imaginary part of complex time (that is the effective coherence time for the particle in units of the decay constant $\Gamma_{2}$), is shown in the lower panel of Fig. 2. The typical range of the time scale $t_{2}$ is between $300~{}\mu s$ and $450~{}\mu s$ (see Fig. 3). Comparing the distribution of the time scale $t_{2}$ to the fugacity spectrum in Fig. 1, and to the time scale $t_{1}$ (upper figure), this shows that the coherent coupling of the single particle quantum states decay on a time scale of about $t_{2}=(300-450)~{}\mu s$. To estimate the time scale in a physical framework, it is assumed that the unit time scale of the defined model is defined by $\tau=2\pi\times(\omega_{x}\omega_{y}\omega_{z})^{-1/3}$, since any frequency of the equilibrated system below that value is effectively zero in the thermally stable quantum system, because of the finite energy uncertainty.

### v6

# Finite phase coherence time of a bosonic quantum field at the Boltzmann equilibrium

@@ Abstract
A quantitative quantum field approach with non-local order parameters is presented for a very weakly interacting, dilute Bose gas. Within the presented model, which assumes the constraint of particle number conservation at constant average energy in the canonical ensemble, it is shown that both coherent oscillations, as well as decay times of quantum coherence for the forward and backward propagating components of the quantum field created by the atomic cloud of a very weakly interacting Bose-Einstein condensate, are defined by a unique time variable. Within the present numerical theory, a quantitative estimate for the unit time scale for time propagation of a particle in a very weakly interacting Bose gas is derived from the coherence time of the wave field and it is illustrated that this time scale defines a unit time for transitions between different realizations of the Boltzmann equilibrium as defined by the maximum entropy from the vanishing of the chemical potential.
Purpose
: arXiv version of the manuscript
@@
As a fundamental property of quantum mechanics, coherence demonstrates the wave nature of quantum particles by the ability of the particles to interfere when their wavelength becomes of the order of their average distance below the critical temperature [2]. In classical physics, when a particle is moved from point A to a second point B, the particle has a well-defined position and velocity at each instant of time during the cycle from A to B, and indeed this time is well-defined by the path length between the two positions and the average velocity of the traveling particle [3]. Quantum mechanically, however, the dynamics of particles are more complicated, because of uncertainty and localization of position and velocity for particles in external confinements, thus in the setup as described above, the quantum particle may only be mainly located at two different positions and therefore exhibit no probability to be located in the space between the two points A and B. As a consequence of quantum effects from quantization of the corresponding single-particle wave functions, if the particle tunnels from point A to point B, there remains a time interval during which it is quantum mechanically not known whether the particle is located at position A or B, so that the quantum state of the particle turns into a coherent superposition of a state with localization in point A and a state with localization in point B, with a mathematically defined phase relation between these two states which allows to quantify the oscillations of the transition probability for the particle to move from point A to B, and vice versa. Apart from spatial localization and coherence as described above, quantum states of non-local type (non-local order parameters) can obey phase coherence, i.e. well-defined relations of the relative or partial phases between the different quantum states. The time scale during which such representing states are quantum mechanically correlated and the particle’s wave function can interfere and coherently oscillate is called coherence time [4].
@@
Calculations and analyses of coherence times typically rely on the knowledge of the spectrum of the given quantum system. In a Bose-Einstein condensate below the critical temperature, most of the particles in the Bose gas search to populate the ground state of the system, so that a macroscopic occupation of the ground state mode with a typically centered spatial distribution occurs below the critical temperature for Bose-Einstein condensation with the characteristics of a coherent matter wave [5]. In the condensed state, all the particles’ wave functions typically tend to constructively interfere at the center of the external confinement. This happens because each of the particles can in principle tunnel from the ground state to a (multimode) excited state of the single-particle spectrum, because of the large energy uncertainty below the critical temperature. As a consequence of such single-particle coherence from tunneling as described above, each particle is thus in general in a superposition of a state with the localization characteristics of the single-particle ground state, and one or multiple of the excited single-particle states - with interfering forward and backward propagating components of the wave fields. The coherence achieved this way is a precursor for Bose-Einstein condensation, where the coherence time is required to be sufficiently large to allow the different single-particle wave functions to interfere constructively and coherently - i. e. in particular larger than the oscillation time of particles among the different single-particle quantum states a particle in the Bose gas may coherently populate.
Within the framework of quantum mechanics, quantum processes obey time-reversal symmetry due to well-defined phase relations, but only on the time scale of coherence, or more precisely, as long as the quantum system is not coupled to an external thermal environment that can quantify the forward propagation of time. It is the oscillation time of forward and backward propagating partial components of the wave fields that defines the time scale for which the two disjoint counterpropagating components interfere at multiples of the oscillation time in complex number representation of the spatially averaged order parameter [6, 7]. As shown in the sequel of this paper, since the oscillation time in complex plane effectively tends to zero on average (standing wave field), it is mainly the coherence time of the quantum field that defines a natural scale for the forward propagation of time. Thus, phase coherence between partial matter waves defines a minimal unit of time in the present many-body quantum system which adjusts its direction of time due to the constructive interference and its scale by the destructive interference of different partial matter waves.
So far, no quantitative estimates for the phase correlation time between counter-propagating partial waves of a quantum particle have been derived from first principles in physical units within standard field theories for Bose-Einstein condensation which make use of particle number conservation in the canonical ensemble. In the following analysis, the coherence time is therefore estimated for a quantum particle in a nearly ideal Bose-Einstein condensate confined in a harmonic trap, within a number-conserving quantum field theory for Bose-Einstein condensation [8, 9], which assumes sufficiently dilute atomic gases. The subsequent theory’s quantum field and its coherence time can be understood as the spatially averaged field with temporal coherence obtained e.g. by the measurement of reflected laser light from the atomic sample.
@@
To connect the mathematical framework of Bose-Einstein condensates at thermal equilibrium to the physics of coherence and oscillation times of a single particle in a Bose gas at finite temperature, the concept of imaginary time turns out to be a suitable mathematical tool to describe both the quantitative scaling of coherence times as well as their uncertainty and broken gauge symmetry aspects induced by many-particle fluctuations below the critical temperature [10]. Whereas the definition of (purely) imaginary time often lacks a clear physical interpretation in other frameworks, such as the calculation of the Hawking radiation in black holes [11], or the definition and calculation of a Eukledian time in special relativity [12], in the present framework, the formal assumption of a complex time variable can be interpreted as the time scale during which a quantum particle in the gas relaxes towards an equilibrium quantum state imposing the assumption of a Boltzmann equilibrium (i. e. coupling elementary quantum states to the external environment).
In comparison to (purely) imaginary time that defines the decay characteristics in the vertical direction of the complex plane, from the conservation of particle number and average energy in the framework of the canonical ensemble, the particles in the gas simultaneously exhibit both finite coherence times as well as oscillations between different regions A and B in space (e. g. spatial domains of different single-particle wave functions) below the critical temperature also at the Boltzmann equilibrium, an effect that can be energetically understood as the coupling of symmetric and asymmetric parts of the underlying complex fugacity spectrum. Thus, the foundation of coherent particle coupling among different condensate and non-condensate atom number states in a Bose-Einstein condensate as described in the present model is due to a process related to elementary time evolution which defines a finite scale for unit time that is larger or equal to the complex time variable with a forward propagating direction as defined by the Boltzmann equlibrium [6]. This in particular defines a (lower bound of) the unit time scale for the forward propagation of time and equivalently the relaxation time for quantum coherence at the interaction of a particle with its (thermal) environment. Therefore, each of the particles is coherently distributed among different single-particle quantum states, which leads to a coherent field distribution with different possible condensate and non-condensate number states at a constant total particle number in the Bose gas, which finally defines a macroscopic ground state population from coherent wave interference at a stable Boltzmann equilibrium of maximum entropy below the critical temperature.
@@
where $\psi_{0}$ represents the spatially integrated condensate field, and $\psi_{\perp}$ the non-condensate field in the framework of solid numbers in the standard complex plane [10]. In the absence of external perturbations (below the critical temperature), and assuming isotropic relative spatial phase dependence, the total field $\psi=0$ vanishes so that the total symmetry of the system remains preserved. The two condensate and non-condensate field components can be decomposed into uncorrelated single field components. We assume a spatially uniform distribution of the relative phase between the different single-particle wave functions. In contrast, as will be shown in the sequel of the present numerical analysis, below the critical temperature, where different atom number states get effectively coherently coupled because of finite coherence times, the resulting quantum states are of the form $|\mu\rangle=\sum_{\bf{k}}c_{\bf{k}}|\{N_{\bf{k}}\}\rangle$ which leads to the fact that $\psi_{0}=-\psi_{\perp}\neq 0$ called spontaneously broken gauge symmetry [13].
Figure 1: (Color online) Figures show the spectrum of fugacity values for a Bose-Einstein condensate with $N=1000$ particles in a harmonic trap with trap frequencies $\omega_{x},\omega_{y}=2\pi\times 42.0$ Hz, and $\omega_{z}=2\pi\times 120.0$ Hz at different temperatures, $T=5.0{\rm~{}nK}$ (upper left), $T=10.0{\rm~{}nK}$ (upper right), $T=22.5{\rm~{}nK}$ (lower left) and $T=35.0{\rm~{}nK}$ (lower right). Below the critical temperature $T_{c}=26.9$ nK, the Boltzmann equilibrium state (Bose-Einstein condensation) corresponds to ${\rm Re}(z)\rightarrow 1$ ($t\rightarrow 0$). In this limit, unconnected parts of the spectrum are strongly coupled. Whereas the fugacity ring spectrum is gapless below the critical temperature, it obeys a gap and tends to get continuously close above $T_{c}$ ($T=35.0{\rm~{}nK}$) with a large range of possible and coupled equilibrium states, obeying the same absolute value of the fugacity tends to one.
Figure 2: (Color online) Figure shows $10^{5}$ realizations of the imaginary parts (upper panel - labeled $t_{1}=\Gamma^{-1}_{1}$) and real parts (lower figure - labeled $t_{2}=\Gamma^{-1}_{2}$) of complex phase as defined in Eq. (3) from random Markov sampling for a non-isotropic trap geometry $\omega_{x},\omega_{y}=2\pi\times 42.0$ Hz and $\omega_{z}=120.0$ Hz at temperature $T=5.0$ nK. The observation that $\Gamma_{2}t$ is non-zero at thermal equilibrium indicates the coupling of the equilibrium state to quantum states with $\Gamma_{2}t\neq 0$ (non-classical correlations).
@@
defines a relation between the absolute temperature and the coherent dynamics of a quantum particle in the presence of the $N-1$ other particles in the Bose gas by its absolute value $|t|=\beta\hbar$. The time parameter $t$ can be interpreted as the absolute value of time for forward time evolution on the resolvable scale of unit time in the equilibrated quantum gas, i. e. the time scale of a stable Boltzmann equilibrium before approaching the next many-particle configuration that corresponds to the equilibrium of maximum entropy. In the present theory, it is the lowest single-particle energy of a quantum particle in the external confinement that defines the uncertainty scale of energy and time, since all values lower than the single-particle ground state energy are effectively zero, because of the finite energy uncertainty.
@@
As a formal ansatz, in Eq. (2), one may use that $\mu t$ is not a real, but a complex number, that is $\mu t\in\mathbb{C}$. Hence, in the framework of this quantum field theory as presented, the real part $\Gamma_{1}t$ describes the coherent phase evolution of the quantum particle at equilibrium (inverse oscillation frequency in the stationary state), corresponding to the ring of the fugacity spectrum in Fig. 1, whereas $\Gamma_{2}t$, the imaginary part (of $\mu t$) describes decay processes of the particle in the atomic cloud at or close to equilibrium (compare linear part of the fugacity spectrum in Fig. 1), that is
$\frac{\mu t}{\hbar\overline{\omega}}=\Gamma_{1}t+i\Gamma_{2}t,$ (3)
where $\overline{\omega}=0.5*(\omega_{x}\omega_{y}\omega_{z})^{1/3}$.
@@
Figure 3: (Color online) Figure shows the coherence phase $t_{2}=\Gamma^{-1}_{2}$ as a function of temperature $T$ for a particle in the trap geometry $\omega_{x},\omega_{y}=2\pi\times 42.0$ Hz and $\omega_{z}=120.0$ Hz. For intermediate values of temperature, the coherence time is around $(300-450)~{}\mu$s. Above the critical temperature, it approaches a value of around $200~{}\mu$s.
Above the critical temperature, the real part of the fugacity spectrum shows a gap between the ring of constant absolute time, indicating that the symmetry of the quantum field can in principle not be broken without externally induced quantum fluctuations of energy and the corresponding entropy. Decreasing temperature by external cooling to decrease the gap between the outer ring and the inner linear (symmetry breaking) part of the fugacity spectrum finally leads to a gapless fugacity spectrum. However, due to the uncertainty of complex time, there is also a finite coupling probability between symmetric and asymmetric parts of the fugacity spectrum, in the parameter range, where the fugacity spectrum obeys a gap between the two principally unconnected spectra close above the critical temperature, which leads to spontaneous symmetry breaking. Thus, assuming Boltzmann equilibrium turns out to be a necessary and sufficient condition for spontaneous symmetry breaking of the quantum field [11] starting from temperatures approaching the critical temperature for Bose-Einstein condensation with a finite coherence time of counter-propagating wave fields (see Fig. 3).
@@
$\Delta N=\sum_{j\neq 0}z^{j}(\mu)\left[\prod_{k=x,y,z}\frac{1}{1-e^{-j\beta\hbar\omega_{k}}}-1\right]+\mathcal{O}(a).$ (5)
In Eq. (5), $\Delta N=(N-N_{0})$ is the difference between $N$ the total number of particles, and $N_{0}$ the number of condensate particles. Further, $\mu$ denotes the chemical potential of the Bose gas, and $\omega_{k}$ the trap frequency in mode direction $k=x,y,z$. For numerical calculations of the conditional probabilities, the single-particle ground state energy (level) $0.5\hbar(\omega_{x}+\omega_{y}+\omega_{z})$ effectively cancels out in the numerical modeling.
Equation (4) does not necessarily describe a coherent forward or backward propagation in time, but only the probability for a transition (or new event) from a state with (the characteristics) of a quantum state with entropy $t_{A}$ to a state with entropy $t_{B}$, that is there is a non-vanishing probability that a transition with $t_{A}\geq t_{B}$ is accepted. Most likely, however, a transition is only accepted, if the absolute value $t_{B}\geq t_{A}$, which means that the particle tends to populate states with large entropy (Boltzmann equilibrium that defines the forward propagation of time) in the Markov evolution process of the present quantum theory. In the present framework, the system is allowed to exchange energy with the environment. i. e. assuming the thermodynamical constraints of the canonical ensemble.
@@
For quantitative numerical modeling and analysis, the Bose gas is assumed to be dilute and obey a constant number of particles at finite temperature in a harmonic trapping potential to numerically calculate the coherence time (real part of complex time as defined in Eq. (2)) of a quantum particle in the Bose gas in SI units, which corresponds to the range of few hundred microseconds at the given parameter space. Interactions in the Bose gas are assumed to be negligible in the limit of a nearly ideal gas, where the s-wave scattering length $a$ of the particles in the gas effectively tends to zero. From Eq. (2), we learn that the Bose gas then obeys a stable (Boltzmann) equilibrium for absolute values of complex time in the limit, where $\Gamma_{1}t\rightarrow 0$ and $|\Gamma_{2}t|\ll\hbar\beta$ (which corresponds to a small decoherence rate as compared to the upper bound $\hbar\beta$ for equilibration in the Bose gas, where all particles populate the single-particle ground state mode in the ideal gas). This indicates that the main decoherence mechanism in the Bose gas is thermalization.
@@
Applying only the constraint of particle number conservation as defined by the conservation equation in Eq. (5), numerical sampling of complex time leads to a distribution of typical time scales for $\Gamma_{1}t$ (oscillation time) and $\Gamma_{2}t$ (coherence time), shown in Fig. 2. From the sampling of complex time, it is observed that $\Gamma_{1}t$, proportional to the real part of complex time, is distributed around zero (with a vanishing width of approximately $0.03~{}\mu s$) in the given parameter range (as defined in Fig. 2 - upper panel). The real part $\Gamma_{1}t$ defines the average oscillation frequency in the fugacity spectrum of Fig. 2, which means that the quantum particle (with constant particle number at finite temperature) quickly couples to excited single particle quantum states, related to $\Gamma_{1}t\sim 0\pm 0.03~{}\mu s$ in the form of a standing wave. The distribution of the time scale $\Gamma_{2}t$, the directional imaginary part of complex time (that is the effective coherence time for the particle weighted with the decay constant $\Gamma_{2}$), is shown in the lower panel of Fig. 2. The typical range of the time scale $\Gamma_{2}t$ is between $300~{}\mu s$ and $450~{}\mu s$ (see Fig. 3). Comparing the distribution of the time scale $\Gamma_{2}t$ to the fugacity spectrum in Fig. 1, and to the time scale $\Gamma_{2}t$ (lower figure), this shows that the coherent coupling of the single particle quantum states decay on a time scale of about $\Gamma_{2}t=(300-450)~{}\mu s$. To estimate the time scale in a physical framework, it is assumed that the unit time scale of the defined model is defined by $\tau=2\pi\times(\omega_{x}\omega_{y}\omega_{z})^{-1/3}$, since any frequency of the equilibrated system below that value is effectively zero in the thermally stable quantum system, because of the finite energy uncertainty.